Solve the following

dy/dx=(x-y-2)/(x+y+6)
1
Expert's answer
2019-09-09T11:41:57-0400

"(x+y+6)dy=(x-y-2)dx"

"(2+y-x)dx+(x+y+6)dy=0"

Let "M=2+y-x, \ N=x+y+6"

"M_y'=1=N_x'" , so the given equation is exact.

And there is function "F(x,y):\ F'_x=M,\ F'_y=N"

"F'_x=2+y-x,\\
 F=\int (2+y-x)dx=2x+yx-x^2/2+h(y)"

"F'_y=(2x+yx-x^2\/2+h(y))'_y=x+h'_y"

"x+h'_y=x+y+6,\ h'_y=y+6,\\
 h(y)=\int (y+6)dy=y^2/2+6y"

"F(x,y)=2x+xy-x^2/2+y^2/2+6y=C,\ C\ is\ constant"

"y^2+2y(6+x)-x^2+4x-2C=0"

Using the quadratic formula

"y=-6-x\pm \sqrt{(6+x)^2-(-x^2+4x-2C)}=\\
-6-x\pm\sqrt{36+12x+x^2+x^2-4x+2C}=\\
-6-x\pm\sqrt{2x^2+8x+36+2C}"

Since "x+y+6" is in the denominator

"x+y+6\not= 0, \ 2x^2+8x+36+2C\not=0.\\
x^2+4x+18+C\not=0\\
(x+2)^2+14+C\not=0\\
x\not=\pm\sqrt{-14-C}-2"

Answer: "y=-6-x\pm\sqrt{2x^2+8x+36+2C}" ,"x\not=\pm\sqrt{-14-C}-2"